# Towards Reliable Human Evaluations in Gesture Generation: Insights from a Community-Driven State-of-the-Art Benchmark

### v2

@@ Abstract
We review human evaluation practices in automated, speech-driven 3D gesture generation and find a lack of standardisation and frequent use of flawed experimental setups. This leads to a situation where it is impossible to know how different methods compare, or what the state of the art is.
@@
Our results provide strong evidence that 1) newer models do not consistently outperform earlier approaches; 2) published claims of high motion realism or speech-gesture alignment may not hold up under rigorous evaluation; and 3) the field must adopt disentangled assessments of motion quality and multimodal alignment for accurate benchmarking in order to make progress.
Finally, in order to drive standardisation and enable new evaluation research, we will release five hours of synthetic motion from the benchmarked models; over 750 rendered video stimuli from the user studies – enabling new evaluations without model reimplementation required – alongside our open-source rendering script, and the 16,000 pairwise human preference votes collected for our benchmark.
@@ 1 Introduction
Research interest in automated gesture generation – the task of animating speaking 3D characters – has been sharply rising as part of the recent boom in generative and multimodal AI [53, 61, 1]. However, whilst the latest generative modelling techniques are being applied to this domain, trustworthy empirical evaluation of communicative non-verbal behaviour remains an understudied problem. Even human evaluations – widely considered the gold standard approach – may lead to misleading conclusions [13, 36] due to their complexity. Therefore, without carefully designed evaluation standards adopted on a community level, published research may give a false sense of progress.
In this paper, we take a closer look at the evaluation practices of recent speech-driven gesture generation works, and identify several dimensions of evaluation setups where common design mistakes may lead to unreliable outcomes, and where the lack of standardisation makes similar studies incompatible. In order to address this problem, we develop an evaluation protocol on BEAT2 [49], the most widely adopted dataset according to our survey. While our protocol is rooted in the GENEA-Challenge methodology, we propose several improvements to increase the validity and re-usability of the evaluations.
To validate our BEAT2 evaluation protocol, and to understand the state of the art in gesture generation, we benchmark six recently published gesture-generation models using crowdsourced human evaluation following the proposed methodology. We then present a discussion of our findings, some of which contradict previously published claims, and release all collected information in the process, in order to enable future research on the evaluation of gesture generation. To summarise, we make three separate contributions to address the lack of standardisation in gesture generation:
@@
    A new evaluation protocol for the widely-used BEAT2 dataset rooted in prior large-scale evaluations (Sec. 4);
@@
    A new benchmark of six recently published models to illuminate the state-of-the-art (Sec. 5).
76.6%10.8%6.8%5.9%
  Not compared
  Mutual baseline (in survey)
  Mutual baseline (not in survey)
  Directly compared
(a) Human evaluation results between pairs of models
77543339
  BEAT   Trinity
  BEAT2   TED
  SHOW   TED-Expr.
  ZEGGS   Other ($\leq 2\times$ used)
(b) Dataset usage
@@ 2.1 Automatic Gesture Generation
We define the task of speech-driven 3D gesture generation as the problem of mapping from a time-aligned input speech sequence $s$ (e.g., an audio waveform and/or a text transcription) with potential additional conditioning $c$ (e.g., an emotion label), to a 3D human body-motion sequence output $x$ (e.g., a series of joint angles defining human poses). Whilst early deep-learning-based systems for gesture generation treated gesture generation as a regression problem [72, 34, 30, 85, 40, 42, 41, 86], recent approaches use probabilistic models like normalising flows [4], VAEs [31], VQ-VAEs [83, 75], discrete autoregressive transformer models [81], or diffusion models [92, 6, 5, 80, 27, 57, 21, 60].
@@ 2.2 Evaluations in Automatic Gesture Generation
Both automated (“objective”) as well as human (“subjective”) evaluation methodologies are widely used in 3D gesture generation [61]. While objective metrics like FGD [86] or beat consistency [52] can be useful to guide model development, they have well-known limitations and unclear relationships with human perception [81, 45, 25, 74, 26, 33, 78]. Therefore, final results are almost universally validated by analysing responses given by human evaluators.
@@ 2.2.1 Standardisation Needs
Despite being considered the gold standard, there is little information available on the ecological validity of human-evaluation practices in gesture generation. A 2021 review by Wolfert et al. [77] found a large diversity in both automated and human evaluation methods, with poor reporting practices on participant characteristics and evaluation design, concluding that the field would “benefit from more experimental rigour and a shared methodology for conducting systematic evaluation”.
The GENEA Challenges [43, 87, 44, 45] were launched in 2020 to address this problem by evaluating sets of gesture-generation models under standardised conditions. The challenges are a series of community-driven evaluations of automated gesture-generation models. Each challenge collected between 5 and 12 model submissions trained on the same data, which were then subjected to large-scale crowdsourced human evaluation by the organisers. Challenge findings emphasise the importance of data filtering, removing artifacts, high-quality visualisations, and disentangled evaluation (as discussed in Sec. 3.1.1), and provide strong overall evidence for the importance of standardised human evaluation.
However, whether the GENEA Challenges address the problem of missing standardisation is an open question. There has been no assessment of whether their methodology is adopted by the community, and their results form isolated user studies rather than a continuously growing benchmark. Furthermore, the GENEA Challenges have seldom included submissions from major computer vision or machine-learning conferences like CVPR or SIGGRAPH, and therefore their results may not reflect the state of the art.
@@ 3 Key Limitations of Current Human-Evaluation Practices
Table 1: Overview of human evaluation practices in 3D gesture-generation research published at SIGGRAPH, SIGGRAPH Asia, and leading computer-vision venues between 2023–2025, as described in Sec. 3. The table uncovers the fragmented state of human evaluation, with inconsistent study designs for related tasks (Tasks column), and a critically low degree of direct comparisons between top models (last column). Abbreviations: SG=SIGGRAPH; Na=Naturalness; Re=Realism, Plausability or Believability; Hu=Human-likeness; Sm=Smoothness; Pref=Preference; Rh=Rhythmic; Sem=Semantic;Gen=General; Em=Emotion; St=Style; B=Present in direct comparison as baseline; and M=Present in direct comparison as main model.
@@
Without standardised benchmarking practices, the state of the art can only be assessed by systematically reviewing independent evaluation results of published models. Are these evaluations reliable, or do inconsistencies in evaluation protocols, participant sampling, and reporting practices undermine their validity? Can we meaningfully assess the state of the art by comparing results from competing models when datasets, visualisation methods, and human-evaluation designs differ so widely across studies, or if such comparisons are fundamentally flawed?
Motivated by these questions, we perform a critical assessment of human evaluation practices in recent gesture-generation research, aiming to answer the following questions about recently published gesture-generation research:
  1. 1.
  2. 2.
We review 26 recent publications on co-speech gesture-generation methods from selected computer vision and graphics conferences (CVPR, ICCV, ECCV, SIGGRAPH, and SIGGRAPH Asia) from 2023 onwards, using the search terms “gesture”, “co-speech”, “speech”, and “motion” in publication titles, then filtering down the results to models whose outputs include 3D body gesture.
BMBMBMBMBMBMBMBMBMBMBMBMBMBMBMBMBMBMBM
@@ 3.1 Entanglement Between Evaluation Dimensions
Upon analysing the human evaluation tasks, we find that gesture-generation research has converged to two high-level goals. In particular, every publication in Tab. 1 tries to assess some variation of the following:
@@
    multimodal alignment between the output motion and the inputs (most commonly, speech).
Each of these aspects is crucial to applications of gesture generation, therefore it is important to assess them independently. A common method is to conduct two similar user studies, one for each aspect, with only the evaluation question changed between the two. We will refer to this as the naive setup.
@@ 3.1.1 The Importance of Mismatching Evaluations
The first GENEA Challenge [43] used the naive setup to evaluate motion quality and multimodal alignment. Surprisingly, models with visually appealing motion received relatively high ratings for multimodal alignment even when they did not depend on the speech. In fact, Kucherenko et al. [45] re-analysed the data for all conditions in the GENEA Challenge 2020 and found a Pearson correlation of over 0.5 between the mean rated appropriateness of a given motion segment and its rated human-likeness. This is despite the fact that their crowdsourced test takers were instructed not to pay attention to human-likeness when assessing appropriateness. Therefore, there is strong evidence that:
Direct evaluations of multimodal alignment – e.g., when conditioning on speech, emotion, style, or semantic information – may be ineffective unless the significant confounding effect of motion realism is isolated.
To combat the above problem, later GENEA Challenges [87, 44, 45] proposed a new evaluation paradigm leveraging the principle of mismatching [29, 68, 38, 87]. In this paradigm, so-called mismatched stimuli are created by swapping the data for a given modality to an unrelated sequence – for example, the audio is replaced by a different sentence, or the movements of the conversational partner are replaced by that of another character. Human evaluators, unaware of the mismatching manipulation, are then asked to indicate their preference between the original (matched) stimulus and its mismatched counterpart, based on which one seems more coherent (e.g., in terms of hand movement matching the speech). This setup effectively isolates the strong confounding effect of intrinsic motion quality from multimodal alignment, as the motion in both stimuli is always generated by the same system. Indeed, the GENEA Challenges 2022 and 2023 confirm the efficacy of this paradigm, reporting distinct model rankings for motion quality versus multimodal alignment [87, 44] and substantially reduced Pearson correlation between the two [45]. However:
Our review finds that none of the multimodal-alignment evaluations in Tab. 1 adopt the mismatching methodology of the GENEA Challenges, nor do they employ other strategies to control for the confounding effect of motion quality.
This methodological gap is critical since it allows models that merely produce smooth, high-quality motion – regardless of how well they align with the speech content – to receive inflated multimodal alignment scores. Consequently, recent findings of speech-gesture alignment in model outputs closely approximating or surpassing the scores of human mocap (e.g., [81, 84, 92, 58]) are not reliable indicators of near-human performance. (We provide strong empirical support for this claim in Sec. 5.4.2).
@@ 3.2 Lack of Direct Comparisons
In this section, we investigate what kind of evaluation results are available when comparing the 26 models included in our survey. The strongest form of evidence for picking between two models comes from a direct comparison. If that is not available, it might still be possible to rank two models if one of them shows larger relative improvement compared to a mutual baseline model. However, if two publications consider non-overlapping sets of models in their evaluation, it becomes infeasible to infer which model works better without running a new evaluation.
We uncover a remarkably low degree of direct comparisons between surveyed models, available only for 19 out of 325 model pairs. As visualised in Fig. 1, this amounts to less than 6% of evaluation coverage of all possible pairs of models in Tab. 1. Shockingly, almost three-quarters of the possible pairings between surveyed publications do not have relevant evaluation results! The right-hand pie chart highlights one reason for this fragmentation: the research community is divided across multiple datasets, with no single benchmark dominating usage. While BEAT and BEAT2 are the most common, a significant proportion of studies use smaller or proprietary datasets, which impedes reproducibility and consistent comparison even further.
It is of course unreasonable to expect every model to be compared to every other model in our review: baselines may be chosen from other conferences, and it is difficult to compare to models trained on other datasets or lacking publicly available implementation. However, human evaluations can depend strongly on which other systems are present in the evaluation, and the range of performance they exhibit [24]. For this reason, the lack of direct comparisons and relevant baseline choices between leading models, as uncovered by our review, pose a serious practical challenge for determining the state of the art.
@@ 3.3 Inconsistent, Incompatible Evaluation Designs
Design choices in, e.g., stimulus creation, question formulation, and test methodology can significantly affect the validity of human evaluation, and may introduce systematic biases. We find that current practices lack standardisation across these dimensions, and therefore it is generally infeasible to draw conclusions from cross-study comparisons.
NaturalHumanlikeSmoothDiverseRealistic0551010101055555544Count
(a) Quality descriptors
SemanticsRhythmUnspecifiedStyle0551010121210106633Count
(b) Forms of alignment
Figure 2: Distribution of what aspects are measured in user studies evaluating motion realism (left) and multimodal alignment (right).
@@
We observe an ad hoc variability in how evaluations formulate the concept of motion realism, with the usage of common adjectives shown on Fig. 2(a). These terms lack a clear definition, and may encode different preferences – for example, “smooth” or “diverse” motion does not necessarily assume that the motion is “human-like” or “believable” While different adjectives could theoretically be used to evaluate distinct styles (e.g., “realism” may be more desirable for photorealistc avatars than for cartoon characters), our survey finds this is not the case in practice. Consequently, we believe that the inconsistency in question formulation needs to be addressed by the community, in order to avoid divergent and incomparable evaluation outcomes.
The lack of standardisation extends to evaluations of multimodal alignment (Fig. 2(b)). While the diversity of evaluation questions can reflect distinct modelling goals – such as semantic grounding, rhythmic alignment, and style control – we have already established that current evaluations setups may fail to isolate such nuances (cf. Sec. 3.1). We therefore postulate that these nominally specialised evaluations in practice measure the same underlying construct of general alignment. This motivates our proposal to establish a standardised, general-purpose alignment task as a foundational measure quantified through mismatching (Sec. 4).
@@
Character visualisation is another critical factor in human evaluation design [55]. Ideally, studies should use high-fidelity 3D avatars that can faithfully reproduce the original motion, avoiding artifacts from retargeting. Reinforcing this, a recent study by Ng et al. [60] found that realistic rendering, compared to untextured meshes, leads to substantially more distinguishing power in evaluations. Despite this, our survey finds that common practice falls short of this ideal. Several recent publications rely on simplistic stick-figure visualisations ([67, 91, 18, 92]), or untextured meshes ([21, 49, 84, 5, 17]). The majority of works in Tab. 1 employ unique, non-standardised 3D characters with substantial differences in realism, artistic style, and even perceived personality traits (Fig. 3). These differences introduce yet another confounding variable to evaluations, undermining the comparability and generalisation of study results.
Response Methodology
@@
Figure 3: Example embodiments used in recent evaluations [92, 81, 87, 62, 21, 60], highlighting the high variety in 3D characters and their degree of realism. There is empirical support for the idea that more expressive visualisations make performance differences substantially easier to spot [60].

### v4

@@ Abstract
We review human evaluation practices in automatic, speech-driven 3D gesture generation and find a lack of standardisation and frequent use of flawed experimental setups. This leads to a situation where it is impossible to know how different methods compare, or what the state of the art is.
@@
Our results show that 1) motion realism has become a saturated evaluation measure on the BEAT2 dataset, with older models performing on par with more recent approaches; 2) previous findings of high speech-gesture alignment do not hold up under rigorous evaluation, even for specialised models; and 3) the field must adopt disentangled assessments of motion quality and multimodal alignment for accurate benchmarking in order to make progress.
To drive standardisation and enable new evaluation research, we release five hours of synthetic motion from the benchmarked models; over 750 rendered video stimuli from the user studies – enabling new evaluations without requiring model reimplementation – alongside our open-source rendering script, and 16,000 pairwise human preference votes collected for our benchmark.
@@ 1 Introduction
Research interest in automatic gesture generation – the task of animating speaking 3D characters – has been sharply rising as part of the recent boom in generative and multimodal AI [53, 59, 1]. However, whilst the latest generative models are being applied to this domain, trustworthy empirical evaluation of communicative non-verbal behaviour remains understudied. Even human evaluations, widely considered the gold standard, may lead to misleading conclusions [11, 35] due to their complexity. Without carefully designed evaluation standards adopted on a community level, published research may give a false sense of progress.
We identify several dimensions of evaluation design in automatic gesture-generation where the lack of standardisation leads to unreliable findings and incomparable studies. To address this, we develop an evaluation protocol on BEAT2 [47], the most widely adopted speech-gesture dataset according to our survey, building on the methodology of the GENEA Challenges [43, 42, 44] with improvements to validity and re-usability.
To validate our protocol, we benchmark six recently published gesture-generation models using crowdsourced human evaluation. We present findings that contradict previously published claims, and release all collected information in the process [^1] Available on https://genea-workshop.github.io/leaderboard/. to enable future research on gesture-generation evaluation. To summarise, our contributions are:
@@
    A new evaluation protocol for the widely-used BEAT2 dataset, rooted in prior large-scale evaluations (Sec. 4);
@@
    A new benchmark of six recent models to illuminate the state of the art (Sec. 5).
@@ 2.1 Automatic Gesture Generation
We define speech-driven 3D gesture generation as the problem of mapping from a time-aligned input speech sequence $s$ (e.g., an audio waveform and/or a text transcription) with potential additional conditioning $c$ (e.g., an emotion label), to a 3D human body-motion sequence output $x$ (e.g., a series of joint angles defining human poses). Whilst early deep-learning-based systems for gesture generation treated gesture generation as a regression problem [70, 33, 29, 84, 39, 40, 41, 83], recent approaches use probabilistic, generative models like normalising flows [5], VAEs [30], VQ-VAEs [81, 73], discrete autoregressive transformer models [79], or diffusion models [90, 7, 6, 78, 25, 57, 19, 58].
@@ 2.2 Evaluations in Automatic Gesture Generation
Both automatic (“objective”) as well as human (“subjective”) evaluation methodologies are widely used in 3D gesture generation [59]. While automatic metrics like FGD [83] or beat consistency [51] can be useful to guide model development, they have well-known limitations and unclear relationships with human perception [79, 44, 23, 72, 24, 32, 76]. Therefore, final results are almost universally validated by analysing responses given by human evaluators.
@@ 2.2.1 Standardisation Needs
Despite being considered the gold standard, there is little information available on the ecological validity of human-evaluation practices in gesture generation. A 2021 review by Wolfert et al. [75] found a large diversity in both automatic and human evaluation methods, with poor reporting practices on participant characteristics and evaluation design, concluding that the field would “benefit from more experimental rigour and a shared methodology for conducting systematic evaluation”.
The GENEA Challenges [42, 85, 43, 44] were launched in 2020 to address this problem by evaluating sets of gesture-generation models under standardised conditions. The challenges are a series of community-driven evaluations of automatic gesture-generation models. Each challenge collected between 5 and 12 model submissions trained on a selected dataset, which were then subjected to large-scale crowdsourced human evaluation by the organisers. Challenge findings emphasise the importance of data filtering, removing artifacts, high-quality visualisations, and disentangled evaluation (as discussed in Sec. 3.1.1), and provide strong evidence for the importance of standardised human evaluation. However, whether the GENEA Challenges address the problem of missing standardisation is an open question. There has been no assessment of whether their methodology is adopted by the community, and their results form isolated user studies rather than a continuously growing benchmark. Furthermore, the GENEA Challenges have seldom included systems from major computer vision or machine-learning conferences like CVPR or SIGGRAPH, and therefore their results may not reflect the state of the art.
@@ 3 Key Limitations of Current Human-Evaluation Practices
We perform a critical assessment of human evaluation practices in recently published gesture-generation research, aiming to answer the following questions:
  1. 1.
  2. 2.
We review 26 recent publications on co-speech gesture-generation methods from selected computer vision and graphics conferences (CVPR, ICCV, ECCV, SIGGRAPH, and SIGGRAPH Asia) from 2023 onwards, using the search terms “gesture”, “co-speech”, “speech”, and “motion” in publication titles, then filtering down the results to models whose outputs include 3D body gesture. The list of papers is presented on Tab. 1.
Table 1: Overview of human evaluation practices in 3D gesture-generation research published at SIGGRAPH, SIGGRAPH Asia, and leading computer-vision venues between 2023–2025, as described in Sec. 3. The table uncovers the fragmented state of human evaluation, with inconsistent study designs for related tasks (Modelling Goal column), and a critically low degree of direct comparisons between top models (last column). Abbreviations: SG=SIGGRAPH; Na=Naturalness; Re=Realism, Plausability or Believability; Hu=Human-likeness; Sm=Smoothness; Pref=Preference; Rh=Rhythmic; Sem=Semantic;Gen=General; Em=Emotion; St=Style; B=Present in direct comparison as baseline; and M=Present in direct comparison as main model.
@@ 3.1 Entanglement Between Evaluation Dimensions
We find a standard practice among reviewed works to conduct human evaluations across two dimensions [^2] See Tab. 1 for details.:
@@
    Multimodal alignment between the output motion and the inputs (most commonly, speech); also known as multimodal grounding [59] and appropriateness [42, 44, 43].
We find that evaluations of motion realism and multimodal alignment are generally conducted using the same user-study setup – where participants compare videos of speaking characters animated by different models – with only the evaluation question changed in between. We will refer to this as the naive approach. Despite its widespread usage, there is evidence that the naive approach fails to separate motion realism and multimodal alignment.
76.6%10.8%6.8%5.9%
  Not compared
  Mutual baseline (in survey)
  Mutual baseline (not in survey)
  Directly compared
(a) Human evaluation results between pairs of models
77543339
  BEAT   Trinity
  BEAT2   TED
  SHOW   TED-Expr.
  ZEGGS   Other ($\leq 2\times$ used)
(b) Dataset usage
@@ 3.1.1 The Importance of Mismatching Evaluations
In 2020, the first GENEA Challenge [42] used the naive approach to evaluate motion realism and speech-gesture alignment. Surprisingly, models with visually appealing motion received relatively high ratings for speech-gesture alignment even when they did not depend on the speech. In fact, Kucherenko et al. [44] re-analysed the data from the GENEA Challenge 2020, and found a Pearson correlation of over 0.5 between the mean rated alignment of a given motion segment and its rated motion realism. This is despite the fact that their crowdsourced test takers were instructed not to pay attention to human-likeness when assessing alignment. Therefore, there is strong evidence that:
Direct evaluations of multimodal alignment – e.g., when conditioning on speech, emotion, style, or semantic information – are ineffective unless one controls for the significant confounding effect of motion realism.
BMBMBMBMBMBMBMBMBMBMBMBMBMBMBMBMBMBMBM
To combat the above problem, later GENEA Challenges [85, 43, 44] proposed a new evaluation paradigm leveraging the principle of mismatching [28, 66, 37, 85]. In this paradigm, so-called mismatched stimuli are created by swapping the data for a given modality to an unrelated sequence – for example, the audio is replaced by a different sentence, or the movements of the conversational partner are replaced by that of another character. Human evaluators, unaware of the mismatching manipulation, are then asked to indicate their preference between the original (matched) stimulus and its mismatched counterpart, based on which one seems more coherent (e.g., in terms of hand movement matching the speech). This setup effectively isolates the strong confounding effect of intrinsic motion quality from multimodal alignment, as the motion in both stimuli is always generated by the same system. Indeed, the GENEA Challenges 2022 and 2023 confirm the efficacy of this paradigm, reporting distinct model rankings for motion quality versus multimodal alignment [85, 43] and substantially reduced Pearson correlation between the two [44].
However, our review finds that none of the multimodal-alignment evaluations in the 26 reviewed works adopt the mismatching methodology of the GENEA Challenges, nor do they employ other strategies to control for the confounding effect of motion quality.
This methodological gap is critical since it allows models that merely produce smooth, high-quality motion – regardless of how well they align with the speech content – to receive inflated multimodal alignment scores. Consequently, recent findings where speech-gesture alignment in model outputs closely approximates or surpasses the scores of human mocap (e.g., [79, 82, 90, 56]) are not reliable indicators of the generated gestures being on par with the dataset. (We provide strong empirical support for this claim in Sec. 5.4.2.)
@@ 3.2 Lack of Direct Comparisons
We investigate what kind of evaluation results are available when comparing the 26 models included in our survey. Unfortunately, we uncover a remarkably low degree of direct comparisons between surveyed models, available only for 19 out of 325 possible model pairs. As visualised in Fig. 1, this amounts to less than 6% of evaluation coverage of all possible pairs of models. Furthermore, almost three-quarters of the possible pairings between surveyed publications do not have relevant evaluation results.
It is unreasonable to expect every model to be compared to every other model in our review: baselines may be chosen from other conferences, and it is difficult to compare to models trained on other datasets or lacking publicly available implementation. At the same time, the lack of direct comparisons between competing models is glaring, since 10 of the 19 comparisons are against weak baseline models introduced with a dataset [82, 47]. Human evaluations can depend strongly on which other systems are present in the evaluation, and the range of performance they exhibit [22]. For this reason, the lack of direct comparisons and relevant baseline choices between leading models, as uncovered by our review, pose a serious practical challenge for determining the state of the art.
0224466881010NaturalHumanlikeSmoothDiverseRealistic
(a) Quality descriptors
0551010SemanticsRhythmUnspecifiedStyle
(b) Forms of alignment
Figure 2: Distribution of aspects measured in user studies evaluating motion realism (left) and multimodal alignment (right).
@@ 3.3 Inconsistent, Incompatible Evaluation Designs
Design choices in, e.g., stimulus creation, question formulation, stimulus presentation, and response format, can significantly affect the validity of human evaluation, and may introduce systematic biases. We find that current practices lack standardisation across these dimensions, and therefore it is generally infeasible to draw conclusions from cross-study comparisons.
@@
We observe an ad hoc variability in how evaluations formulate the concept of motion realism, with the usage of common adjectives shown on Fig. 2(a). These terms lack a clear definition, and may encode different preferences – for example, “smooth” or “diverse” motion does not necessarily assume that the motion is “human-like” or “believable”. While different adjectives could theoretically be used to evaluate distinct styles (e.g., “realism” may be more desirable for photorealistic avatars than for cartoon characters), our survey finds this is not the case in practice. The lack of standardisation extends to evaluations of multimodal alignment (Fig. 2(b)). Most evaluations attempt to isolate specific forms of alignment: expression of meaning and of style, or temporal alignment between speech and gesture, but we also find six publications evaluating general (unspecified) alignment to the speech. While the diversity of evaluation questions can reflect distinct modelling goals, the employed naive evaluation setups may fail to isolate such nuances (cf. Sec. 3.1).
@@
Character visualisation is another critical factor in human evaluation design [54]. Ideally, studies should use high-fidelity 3D avatars that can faithfully reproduce the original motion, avoiding artifacts from retargeting. Reinforcing this, a recent study by Ng et al. [58] found that realistic rendering, compared to untextured meshes, leads to substantially more distinguishing power in evaluations. Despite this, our survey finds that common practice falls short of this ideal. Several recent publications rely on simplistic stick-figure visualisations ([65, 89, 16, 90]), or untextured meshes ([19, 47, 82, 6, 15]). The majority of surveyed evaluations employ unique, non-standardised 3D characters with substantial differences in realism, artistic style, and even perceived personality traits (Fig. 3). These differences introduce yet another confounding variable to evaluations, undermining the comparability and generalisation of study results.
User-Study Response Format
@@
Figure 3: Example embodiments used in recent evaluations [90, 79, 85, 60, 19, 58], highlighting the high variety in 3D characters and their degree of realism. There is empirical support for the idea that more expressive visualisations make performance differences substantially easier to spot [58, 26].